OpenStudy (anonymous):

Is the sequence 5, 9, 15, ... an arithmetic sequence? Explain.

OpenStudy (experimentx):

no ... arithmetic sequence should be like this 5,9,13,17, ... constant difference

OpenStudy (anonymous):

i not sure to to explain that its not

OpenStudy (experimentx):

5,9,13,17, 9 = 5+4 13 = 9+4 17 = 13+4

OpenStudy (anonymous):

so how do i explain it in words and i have no i dea what any of that is :(

OpenStudy (mertsj):

It is arithmetic if you always add the same number to get the next term.

OpenStudy (mertsj):

Just say "It is not arithmetic because there is no common difference."

OpenStudy (anonymous):

ok thank you very much

OpenStudy (mertsj):

yw